OpenStudy (lexy148):

how would you solve this to put it into y=mx+b form. 2x+8=0

OpenStudy (anonymous):

Assuming the general linear function that describes 2x+8: y = 2x+8 when y = 0: 2x+8 = 0

OpenStudy (lexy148):

ok